OpenStudy (yaboy):

Choose the equation below that represents the line that passes through the point (−2, −1) and has a slope of 5. A.) y − 1 = 5(x − 2) B.) y + 1 = 5(x + 2) C.)y + 2 = 5(x + 1) D.) y − 2 = 5(x − 1)

OpenStudy (eliesaab):

The equation of line going through a point (a,b) and has slope m is y-b=m(x-a)

OpenStudy (eliesaab):

You can now find your choice

OpenStudy (yaboy):

so a or d?

OpenStudy (eliesaab):

No if a =-2, then (x-a)=(x+2)

OpenStudy (eliesaab):

If b=-1, then x-b=y+1